OpenStudy (anonymous):

Part 1: Write the equation of the line that passes through the points (–1, –4) and (2, 5). Part 2: Using complete sentences, explain whether or not it matters which point is used in the final answer. Also explain why you chose the point you did.

hero (hero):

Find the slope, then use point slope formula. It won't matter which point you choose

OpenStudy (anonymous):

Psssssssssst Follow Hero...

OpenStudy (anonymous):

You're annoying -.-

OpenStudy (anonymous):

Lol sorry but If you need help you can ask me psssst

OpenStudy (anonymous):

What is the point slope formula?

hero (hero):

m(y - y_1) = (x - x_1)

OpenStudy (anonymous):

\[m=\frac{y_2-y_1}{x_2-x_1}\]

hero (hero):

That's the slope formula, not the point-slope formula

OpenStudy (anonymous):

Right and i got m = 3

OpenStudy (anonymous):

So what now?

OpenStudy (anonymous):

Follow Hero

hero (hero):

3(y - 5) = (x - 2)

hero (hero):

There ya go

OpenStudy (anonymous):

And that's it?

hero (hero):

Nope, you have to simplify it into the form y = mx + b

OpenStudy (anonymous):

Yup that's it

OpenStudy (anonymous):

I mean a little simplification more

OpenStudy (anonymous):

and then you're done

OpenStudy (anonymous):

d-_-b

OpenStudy (anonymous):

I'm still confused. What's b?

OpenStudy (anonymous):

b constant

hero (hero):

Simplify it first and isolate y

hero (hero):

I'll let you know what b is later

OpenStudy (anonymous):

3y - 15 =x +2 ??

OpenStudy (anonymous):

Did i do that wrong?

hero (hero):

How'd you get +2 ?

OpenStudy (anonymous):

My bad. 3y - 15 = x - 2

hero (hero):

good, now keep simplifying until you get y by itself

OpenStudy (anonymous):

What comes next though? subtract 2?

hero (hero):

No.....add 15 to both sides

OpenStudy (anonymous):

So.. 3y = x + 13?

hero (hero):

Yes, and now divide both sides by 3

OpenStudy (anonymous):

y = 3x +13 over 3?

hero (hero):

\[y = \frac{x}{3} + \frac{13}{3}\]

OpenStudy (anonymous):

Right

OpenStudy (anonymous):

You ^^^^^^^^^ have it

OpenStudy (anonymous):

Okay so now what?

OpenStudy (anonymous):

Lol stop

OpenStudy (anonymous):

that's it

hero (hero):

b = 13/3

OpenStudy (anonymous):

*Dead End*

OpenStudy (anonymous):

So then what?

hero (hero):

That's it

hero (hero):

We're done

hero (hero):

Now go celebrate with raw cashews and cinna honey

OpenStudy (anonymous):

No what's the final equation though? lmao